A firm has an inventory turnover rate of 10, a receivables turnover rate of 14, and a payables turnover rate of 12. How long is the operating cycle?

71.80 days

68.72 days

44.79 days

62.57 days

54.15 days
Answer #1

Calculation of operating cycle:

Operating cycle= days sale of inventory + days sale outstanding

Days sale of inventory= 365/inventory turnover= 365/10= 36.5days

Days sale outstanding= 365/recievables turnover rate= 365/14= 26.07days

Operating cycle= 36.5+26.07= 62.57days

So correct answer is 62.57 days